2x+3y=10 , find the value of x and y.
Answer by confusedgirl81(14)

with this you pick your own "x" or "y" so,
x=-1
2x+3y =10
2(-1)+3y =10 substitute -1 for x
-2+3y =10 add 2 to both sides
3y =12 divide by 3
y =4 (-1,4)

x=2
2x+3y =10
2(2)+3y =10 substitute 2 for x
4+3y =10 subtract 4 from both sides
3y =6 divide by 3
y =2 (2,2)

choose any x then solve for y.